OpenStudy (anonymous):

find the distance from (-3,2) to (5,-13)

OpenStudy (anonymous):

\[(x_2-x_1)/(y_2-y_1)\] = (5- -3)/(-13-2) =8/15 ... I could be wrong though - haven't done this in a while. someone else should check :)